OpenStudy (anonymous):

verify my toyo... please (toyo=mess) integrate sin^5x cos^2x dx

OpenStudy (anonymous):

=-1/3cos^3x + 2/5 cos^5x - 1/7cos^7x +c

OpenStudy (anonymous):

i used easy sub..

OpenStudy (anonymous):

make u = cos(x), du -sin(x) dx then integrate \[-\int u^2(1-u^2)^2 du\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sometimes these are a pain, i think this one is ok yes?

OpenStudy (anonymous):

so i am rigth? right

OpenStudy (anonymous):

i didn't do it but i bet it is. do you integrate what i wrote?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it is right!